OpenStudy (anonymous):

Factor: 4v^2-16v+12=0

OpenStudy (anonymous):

(v-3)(v-1)

OpenStudy (xapproachesinfinity):

irrelevant! you should not give answers directly! you have to explain how you achieved the answer! best leave the asker see through the whole problem!

OpenStudy (xapproachesinfinity):

Notice that 4 is GCF btw all the terms?

OpenStudy (xapproachesinfinity):

factor that out

OpenStudy (xapproachesinfinity):

\(\large 4(v^2-4v+3)=0\) surely 4 cannot be zero then is must be \(\large v^2-4v+3=0\) now factor this we ask what are the two numbers that multiply to get 3 and add up to -4

OpenStudy (xapproachesinfinity):

to get that we look at factor of 3 {1,3} 1 and 3 multiply to 3 but the add up to 4 not -4 so we need to put some negative sign to them since we have positive 3 then we should have -1 and -3 adding them we get -4 so \(\large (v-3)(v-1)=0\)

OpenStudy (xapproachesinfinity):

no solve for v